# A flexible, GPU-accelerated approach for the joint characterization of LISA instrumental noise and Stochastic Gravitational Wave Backgrounds

### v1

@@ III.4 Two-step inference
Our last acceleration step consists of splitting the inference into two steps. There is a degeneracy in the model in that changes to the model template parameters can be mimicked by changes to the spline parameters. We address this degeneracy by first running an initial MCMC with only the modelled noise parameters and the background templates. This corresponds to enforcing all the spline knots to have $w_{k}=0$ in Eq. (14). The resulting posterior distributions are used to produce the results labelled as “template recovery” in Figs. 1 and 4, and to fix the template parameters in a second MCMC run. In the second run, we let the splines free to explore the parameter space, while the template parameters $\boldsymbol{\lambda}$ are fixed to the medians of the aforementioned posteriors. If compared to running one single MCMC where all the parameters $\boldsymbol{\theta}$ are updated together, this separation dramatically improves the convergence of the chains, as we remove the degeneracy between the template parameters and the splines. We can do this because in our model the splines are not used to describe the overall spectral shape of the PSD, but only to account for (small) deviations from the templates. We can interpret the recovered template model parameters as the best fit model, and the spline parameters as deviations from that best fit. This applies to both the noise and the signal analyzed. In general, if the spline model was used to describe the totality of the stochastic contribution as one individual component, the priors would be increased to be as large as needed. Instead, we interpret the splines as deviations from the best-fit model; therefore, we use tight zero-mean Gaussians as priors on $w_{k}$.
@@ IV.1 Cosmic strings
[…]
@@
The situation is different for the splines analysis in Fig. LABEL:fig:_cs_splines. Under the first hypothesis, the reconstructed $90\%$ credible regions, represented as filled contours, entirely encompass the injected PSDs. In the second, the reconstruction perfectly overlays with the averaged data periodogram shown in grey. In both panels of Fig. 1, the $90\%$ credible regions for the noise recovery under both assumptions are tight enough to be indistinguishable from the median of the posterior distributions. We can compute the Bayes factor between the two hypotheses to assess which is better supported by the data. This is a pivotal question posed by resorting to flexible methods, as lifting the assumption of knowing the spectral shape of the different contributions inevitably reduces our constraining power. Flexible noise models can accommodate the excess power present in the data because of a stochastic signal, making it harder to detect an SGWB and constrain its parameters. Labeling the first hypothesis as “noise+signal” ($ns$) and the second as “noise” ($n$), we compute the respective evidences using Eq. (16) and the natural logarithm of the Bayes factor $\log\mathcal{B}_{ns/n}=\log Z_{ns}-\log Z_{n}$. We find a log-Bayes factor of $\log\mathcal{B}_{ns/n}=6.54\pm 0.44$, which, according to the criterion outlined in Ref. [103], translates to “very strong” evidence for the presence of a stochastic signal in the data. Thus, in this case, our flexible parametrization allows us to confidently discriminate between the presence or absence of an SGWB. Although the splines can absorb the SGWB, they do so by varying considerably more parameters, so the more parsimonious description of an SGWB is still preferred. Nevertheless, we want to stress that the introduction of the splines reduced the log-Bayes factor by three orders of magnitude, as shown in Fig. 1. Therefore, the increased flexibility introduced in the noise (and signal) model(s) could make detecting SGWBs with SNR lower than the one considered much more challenging, as also shown in Ref. [43]. This is particularly relevant for SGWBs that are quieter than our example, but that would still be detectable in principle in a templates-only analyses.
@@
The interpretation of the absolute value of the Bayes factor in the context of SGWB analyses is an active area of research with the authors of Refs. [43, 41, 42, 44]. Nevertheless, here it still provides a valuable metric to assess the impact of model-agnostic analyses on the detectability of these signals. We also refer the reader to [104], where the impact of the inclusion of model uncertainties in the Bayes factor computation was investigated.
@@
Given the stochastic nature of the datasets, we would expect that flexible models are capable of achieving larger likelihoods than the analytic templates, even if the latter ones match the spectral shape used during the data generation. This is a consequence of the flexible models being able to better accommodate the random fluctuations inherent to stochastic processes. Both the noise and signal datasets simulated here are random realizations whose PSD is described by Eqs. (6) and (9) only on average. Therefore, we expect the presence of local fluctuations that deviate from the templates. Given the local nature of the splines, they can account for these fluctuations, while the templates cannot. In principle, one could reach larger and larger likelihood values by adding knots until the recovered PSD exactly matches the data periodogram everywhere. This effect is balanced by the increased prior volume associated with the additional parameters. The fact that only a small number of interior knots are added when the data is generated from a PSD consistent with the template family gives us confidence that our priors on the spline parameters have been chosen well.
@@
We report the posterior on the number of in-between spline knots in the $A$ channel in Fig. 2. The posteriors in $A$ and $E$ always look the same, since the PSD in the two channels has the same shape. Conversely, given its low sensitivity to GWs, the $T$ channel posteriors always converge to one single knot. Colors reflect the labels already described for Fig. 1. Under the “noise+signal” hypothesis, all the spline perturbations converge to having only one in-between knot. On the other hand, given the loud nature of the signal, the recovery under the noise-only hypothesis requires between a total of 5 and 6 knots to account for the totality of the power present in the data stream. This behavior reflects our expectations. Since the simulated data are consistent with the templates used during the inference stage, we do not have to rely on the spline component to catch deviations from the parametrized baseline. Nevertheless, since one extra knot is still added by the reversible jump algorithm, we believe that its role is to account for the aforementioned random fluctuations. The different posteriors help interpret the Bayes factors previously presented. In fact, even if our flexible model under the “noise-only” hypothesis can provide a good description of the data, it requires a larger number of parameters than the model including both noise and signal. The Bayes factor inherently includes, in some sense, the concept of the “Occam’s razor”, i.e., if two models describe the data equally well, the more complex one is disfavoured. The Bayes factors found here demonstrate that we prefer the simplest description of the data in this case, which is the noise model template plus a SGWB.
@@
Figure 1 shows the $90\%$ credible regions getting wider when splines are included in the model. The effect is more pronounced for the SGWB contribution, while it is barely visible for the remaining components. To better highlight this behaviour, we show in Fig. 3 the posterior distribution of the reconstructed total PSD in the three TDI channels at a reference frequency of $\bar{f}=10^{-3}\,\rm Hz$. We note that the colorscheme in this Figure is different to the one used in Figs. 1 and 2. This Figure allows us to compare how the location and width of the PSD posteriors change with the assumed model. The tightest contours are obtained by the two template only analysis. The noise-only recovery in purple is clearly biased in all three channels. Including the splines in the model has the effect of broadening the posteriors for both the “noise+signal” and the “noise” recoveries, and to shift the latter enough to be in agreement with the injection.
@@ IV.2 Perturbed noise spectrum
[…]
@@
[…]
@@
The library used to perform the analysis presented in this paper is already available at https://github.com/asantini29/lisa-ps. This package is still under early development, and it is likely to undergo significant changes in the future. We also release our parallel, GPU-accelerated implementation of Akima splines at https://github.com/asantini29/CudAkima [92]. Besides the few packages already cited, this work made extensive use of Numpy [105], Matplotlib [106], Numba [107], and Pysco [108].

### v2

@@ III.4 Two-step inference
Our last acceleration step consists of splitting the inference into two consecutive steps. There is a degeneracy in the model in that changes to the model template parameters can be mimicked by changes to the spline parameters. We address this degeneracy by first running an initial MCMC with only the modelled noise parameters and the background templates. This corresponds to enforcing all the spline knots to have $w_{k}=0$ in Eq. (14). The resulting posterior distributions are used to produce the results labelled as “template recovery” in Figs. 1 and 4, and to fix the template parameters in a second MCMC run. In the second run, we let the splines free to explore the parameter space, while the template parameters $\boldsymbol{\lambda}$ are fixed to the medians of the aforementioned posteriors. If compared to running one single MCMC where all the parameters $\boldsymbol{\theta}$ are updated together, this separation dramatically improves the convergence of the chains, as we remove the degeneracy between the template parameters and the splines. We can do this because in our model the splines are not used to describe the overall spectral shape of the PSD, but only to account for (small) deviations from the templates. We can interpret the recovered template model parameters as the best fit model, and the spline parameters as deviations from that best fit. This applies to both the noise and the signal analyzed. In general, if the spline model was used to describe the totality of the stochastic contribution as one individual component, the priors would be increased to be as large as needed. Instead, we interpret the splines as deviations from the best-fit model; therefore, we use tight zero-mean Gaussians as priors on $w_{k}$.
@@
This two-steps approach, combined with the hardware acceleration and averaging procedure described in Secs. III.2 and III.3, allows our runs to converge and accumulate enough samples within a day on an NVIDIA A100-SXM4 GPU despite the large number of parallel walkers and the dimensionality of the parameter space.
@@ IV.1 Cosmic strings
Figure 1: Reconstructed posterior distribution of the different contributions to the total PSD in the AA channel. The black dashed (dash-dotted) line represents the PSD used to generate the noise (SGWB from cosmic strings) dataset. The averaged data periodogram is shown as a grey line. Blue (orange) lines and filled contours represent medians and $90\,\%$ credible regions for the reconstructed noise (signal) PSD when both contributions are present in the model. Red lines and contours refer to the case where only the noise has been used to fit the totality of the data. Insets show a zoom-in on the region $f\in[2.5,\,4]\,\rm mHz$. Left panel: Result of the analysis performed using only the baseline templates Right panel: Result of the analysis performed introducing splines in the model.
@@
The situation is different for the splines analysis in Fig. LABEL:fig:_cs_splines. Under the first hypothesis, the reconstructed $90\%$ credible regions, represented as filled contours, entirely encompass the injected PSDs. In the second, the reconstruction perfectly overlays with the averaged data periodogram shown in grey. In both panels of Fig. 1, the $90\%$ credible regions for the noise recovery under both assumptions are tight enough to be mostly indistinguishable from the median of the posterior distributions. We can compute the Bayes factor between the two hypotheses to assess which is better supported by the data. This is a pivotal question posed by resorting to flexible methods, as lifting the assumption of knowing the spectral shape of the different contributions inevitably reduces our constraining power. Flexible noise models can accommodate the excess power present in the data because of a stochastic signal, making it harder to detect an SGWB and constrain its parameters. Labeling the first hypothesis as “noise+signal” ($ns$) and the second as “noise” ($n$), we compute the respective evidences using Eq. (16) and the natural logarithm of the Bayes factor $\log\mathcal{B}_{ns/n}=\log Z_{ns}-\log Z_{n}$. We find a log-Bayes factor of $\log\mathcal{B}_{ns/n}=6.54\pm 0.44$, which, according to the criterion outlined in Ref. [109], translates to “very strong” evidence for the presence of a stochastic signal in the data. Thus, in this case, our flexible parametrization allows us to confidently discriminate between the presence or absence of an SGWB. Although the splines can absorb the SGWB, they do so by varying considerably more parameters, so the more parsimonious description of an SGWB is still preferred. Nevertheless, we want to stress that the introduction of the splines reduced the log-Bayes factor by three orders of magnitude, as shown in Fig. 1. Therefore, the increased flexibility introduced in the noise (and signal) model(s) could make detecting SGWBs with SNR lower than the one considered much more challenging, as also shown in Ref. [45]. This is particularly relevant for SGWBs that are quieter than our example, but that would still be detectable in principle in a templates-only analyses.
@@
The interpretation of the absolute value of the Bayes factor in the context of SGWB analyses is an active area of research with the authors of Refs. [45, 43, 44, 46]. Moreover, it is always influenced by the prior distributions assumed during the analysis. Nevertheless, here it still provides a valuable metric to assess the impact of model-agnostic analyses on the detectability of these signals. We also refer the reader to Ref. [110], where the impact of the inclusion of model uncertainties in the Bayes factor computation was investigated. For comparison, we also apply the Deviance Information Criterion (DIC) [111, 112], an alternative model selection tool that does not require the computation of the evidence. We define the deviance as $D(\boldsymbol{\theta})=-2\log\mathcal{L}(d|\,\boldsymbol{\theta})+C$, where $C$ is a constant that cancels out in all model comparisons. Using the notation $\overline{x}$ to indicate the mean over the posterior distribution of the quantity $x$, we can define the effective number of model parameters as $p_{D}=\overline{D(\boldsymbol{\theta})}-D(\overline{\boldsymbol{\theta}})$. From these quantities, the DIC is computed as
$\text{DIC}=\overline{D(\boldsymbol{\theta})}\,+\,p_{D}.$ (20)
Models with smaller DIC values should be preferred. In particular, Ref. [109] suggests that the absolute value of the difference in DIC between two models, $|\Delta\text{DIC}|$, should roughly follow the same scale of twice the natural logarithm of the Bayes. For the template analysis, we find $\Delta\text{DIC}=\text{DIC}_{ns}-\text{DIC}_{n}=-1.3\times 10^{4}$, while in the splines analysis we obtain $\Delta\text{DIC}=-30$. This agrees with our Bayes factor results, favouring the “noise+signal” hypothesis in both cases. Once again, the inclusion of splines significantly reduces the strength of the support for the presence of a signal. However, when analysing the “noise+signal” data, we find $\text{DIC}_{ns}$ to be (slightly) smaller in the templates-only analysis, while the splines case shows a (slightly) larger evidence than the templates-only case. Therefore, under our inference setup, the DIC seems (slightly) more penalizing towards increased model complexity than the Bayes factor.
@@
Given the stochastic nature of the datasets, we would expect that flexible models are capable of achieving larger likelihoods than the analytic templates, even if the latter ones match the spectral shape used during the data generation. This is a consequence of the flexible models being able to better accommodate the random fluctuations inherent to stochastic processes. Both the noise and signal datasets simulated here are random realizations whose PSD is described by Eqs. (6) and (9) only on average. Therefore, we expect the presence of local fluctuations that deviate from the templates. Given the local nature of the splines, they can account for these fluctuations, while the templates cannot. In principle, without an upper bound on the number of knots, one could reach larger and larger likelihood values by adding knots until the recovered PSD exactly matches the data periodogram everywhere. This effect is balanced by the increased prior volume associated with the additional parameters. The fact that only a small number of interior knots are added when the data is generated from a PSD consistent with the template family gives us confidence that our priors on the spline parameters have been chosen well.
@@
We report the posterior on the number of in-between spline knots in the $A$ channel in Fig. 2. The posteriors in $A$ and $E$ always look the same, since the PSD in the two channels has the same shape. Conversely, given its low sensitivity to GWs, the $T$ channel posteriors always converge to one single knot. Colors reflect the labels already described for Fig. 1. Under the “noise+signal” hypothesis, all the spline perturbations converge to having only one in-between knot. On the other hand, given the loud nature of the signal, the recovery under the noise-only hypothesis requires between a total of 5 and 6 knots to account for the totality of the power present in the data stream. This behavior reflects our expectations. Since the simulated data are consistent with the templates used during the inference stage, we do not have to rely on the spline component to catch deviations from the parametrized baseline. Nevertheless, since one extra knot is still added by the reversible jump algorithm, we believe that its role is to account for the aforementioned random fluctuations. The different posteriors help interpret the Bayes factors previously presented. In fact, even if our flexible model under the “noise-only” hypothesis can provide a good description of the data, it requires a larger number of parameters than the model including both noise and signal. Usually, we would like our analyses to implement the concept of “Occam’s razor,” i.e., if two models describe the data equally well, the more complex one is disfavoured. It’s often said that the Bayes always automatically includes it, but this holds only in cases where suitable priors are chosen (see, e.g., Ref [113]). The Bayes factors found here demonstrate that our priors are chosen appropriately, since we prefer the simplest description of the data in this case, which is the noise model template plus an SGWB.
@@
Figure 1 shows the $90\%$ credible regions getting wider when splines are included in the model. The effect is more pronounced for the SGWB contribution, while it is barely visible for the remaining components. To better highlight this behaviour, we show in Fig. 3 the posterior distribution of the reconstructed total PSD in the three TDI channels at a reference frequency of $\bar{f}=10^{-3}\,\rm Hz$. We note that the colorscheme in this Figure is different to the one used in Figs. 1 and 2. This Figure allows us to compare how the location and width of the PSD posteriors change with the assumed model. The tightest contours are obtained by the two template only analysis. The noise-only recovery in purple is clearly biased in all three channels. Including the splines in the model has the effect of broadening the posteriors for both the “noise+signal” and the “noise” recoveries, and to shift the latter enough to be in agreement with the injection. Both here and in Fig. LABEL:fig:_cs_splines, the credible intervals for the splines analysis account only for the uncertainty in the spline parameters, while the template parameters remain fixed. The same applies to the evidence and DIC calculations in the splines analyses. While being a caveat, the posteriors in Fig. 3 suggest that the uncertainty in the spline coefficients translates in larger contours than the uncertainty in the template parameters alone. Therefore, we do not expect our conclusions to change significantly.
@@ IV.2 Perturbed noise spectrum
Finally, we present our results for the “perturbed” noise PSD. In this case, we have generated the data according to a PSD of the same shape as Eq. (14a) with $\mathcal{P}_{n,ii}(f)$ being splines with knot positions $\log_{10}f_{k}=\left\{-7.5,\,-4,\,-2.01,\,-1.71,\,-1.54,\,-0.6\right\}$ and amplitudes $w_{k}=\left\{0.0,\,0.2,\,-0.1,\,0.3,\,0.2,\,0.0\right\}$. We assume the same knot positions and amplitudes for all the channels, and we include the leftmost and rightmost points to cover the whole frequency range implied by our assumed observation period and cadence. The base, unperturbed noise PSD and the one used for the injection are shown as the black dash-dotted and dashed lines in Fig. 4. Red and blue lines and contours represent the median and $90\%$ credible interval for the reconstructed PSD for the template and spline recoveries, respectively. While the first approach fails to successfully represent the power present in the data, the spline analysis is able to recover the injected PSD despite it being different from the assumed baseline. This gives us confidence that the results of the previous section do not strongly depend on having the right template as the baseline of the flexible models, as long as the differences in the spectral shape are small enough to be compatible with our priors.
@@
Figure 4: Reconstructed posterior distribution of the reconstruction of the total PSD in the AA channel. The black dashed (dash-dotted) line represents the “perturbed” PSD used to generate the noise dataset (the base noise PSD before applying the spline perturbation). The averaged data periodogram is shown as a grey line. Blue lines and filled contours represent the median and $90\%$ credible interval for the reconstructed noise PSD. The inset shows a zoom-in on the region $f\in[3.1,\,4]\,\rm mHz$.
@@
The library used to perform the analysis presented in this paper is already available at https://github.com/asantini29/lisa-ps [114]. This package is still under early development, and it is likely to undergo significant changes in the future. We also release our parallel, GPU-accelerated implementation of Akima splines at https://github.com/asantini29/CudAkima [96]. Besides the few packages already cited, this work made extensive use of Numpy [115], Matplotlib [116], Numba [117], and Pysco [118].
@@ Acknowledgements.
Appendix A Akima splines implementation
Given a set of knot points with coordinates ${(x_{i},\,y_{i})}_{i=1,\dots,n}$, Akima splines construct a piecewise function $S(x)$ such that $S(x_{i})=y_{i}$ for all $i$. In those points, its slope $s_{i}$ is a function of $(x_{j},\,y_{j})$ with $j\in\{i-2,\dots,i+2\}$. We can define the slope of the line connecting two consecutive knots $(x_{i},\,y_{i})$ and $(x_{i+1},\,y_{i+1})$ as
$m_{i}=\frac{y_{i+1}-y_{i}}{x_{i+1}-x_{i}},\quad i=0,\dots,\,n-1.$ (21)
Defining the quantity
$w_{i}=|m_{i+1}-m_{i}|,$ (22)
we can construct the spline slopes as
$s_{i}=\begin{dcases}\frac{m_{i+1}+m_{i}}{2}\quad&\text{if }w_{i}+w_{i-2}=0,\\ \frac{w_{i}\,m_{i-1}+w_{i-2}\,m_{i}}{w_{i}+w_{i-2}}\quad&\text{otherwise}.\end{dcases}$ (23)
For each knot $(x_{i},y_{i})$ the coefficients $s_{i}$ require 4 extra points to be computed, therefore we need to introduce boundary conditions for the first and last two knots. In our implementation [96], we follow Ref. [88]:
$s_{1}$ $=\frac{3m_{1}-m_{2}}{2};$ (24a)
$s_{2}$ $=\begin{dcases}\frac{m_{3}+m_{2}}{2}\quad&\text{if }w_{2}+w_{1}=0\\ \frac{w_{2}\,m_{1}+w_{1}\,m_{2}}{w_{2}+w_{1}}&\text{otherwise};\end{dcases}$ (24b)
$s_{n-1}$ $=\begin{dcases}\frac{m_{n-3}+m_{n-2}}{2}\quad\text{if }w_{n-3}+w_{n-2}=0\\ \frac{w_{n-3}\,m_{n-2}+w_{n-2}\,m_{n-3}}{w_{n-3}+w_{n-2}}\quad\text{otherwise};\end{dcases}$ (24c)
$s_{n}$ $=\frac{3m_{n-1}-m_{n-2}}{2}.$ (24d)
Since the spline is defined as a piecewise cubic polynomial, in each interval $x\in[x_{i},\,x_{i+1})$ we have:
$S(x)=a_{i}+b_{i}(x-x_{i})+c_{i}(x-x_{i})^{2}+d_{i}(x-x_{i})^{3}.$ (25)
The coefficients are determined by imposing the continuity of $S(x)$ and its first derivative $S^{\prime}(x)$ at each knot:
$S(x_{i})=y_{i},$ $\quad S(x_{i+1})=y_{i+1}$ (26a)
$S^{\prime}(x_{i})=s_{i},$ $\quad S^{\prime}(x_{i+1})=s_{i+1}.$ (26b)
Therefore, the polynomial coefficients read:
$a_{i}$ $=y_{i},$ (27a)
$b_{i}$ $=s_{i},$ (27b)
$c_{i}$ $=\frac{3m_{i}-2s_{i}-s_{i+1}}{x_{i+1}-x_{i}},$ (27c)
$d_{i}$ $=\frac{s_{i}+s_{i+1}-2m_{i}}{(x_{i+1}-x_{i})^{2}}.$ (27d)
Since there are no conditions on the second derivative of $S(x)$, this is not necessarily continuous.